OpenStudy (anonymous):

A body falls freely from rest. It covers as much distance in the last second of its motion as covered in first three seconds. Calculate the total time of fall for the body.

OpenStudy (anonymous):

@dan815 @Compassionate @Nnesha @Kainui

OpenStudy (shamim):

For first 3sec S=0.5a*3^2 S=4.5a

OpenStudy (shamim):

For last sec S=0.5a*(2t-1)

OpenStudy (shamim):

According to question 4.5a=0.5a(2t-1) 4.5=0.5(2t-1) t=?

OpenStudy (shamim):

Response plz!!!!!

OpenStudy (shamim):

We can write a=g too